Chlorine gas can be produced by the reaction of hydrochloric acid with manganese(IV) oxide. If the percent yield for the reaction is 91%, how many grams of chlorine will be produced by the reaction of 212 g of HCl with excess MnO2?
1
Expert's answer
2020-05-19T08:47:03-0400

"MnO_2(s) + 4HCl(aq) \u2192 Cl_2(g) + MnCl_2(aq) + 2H_2O(l)"

from the above equation we can see that for every 4 mole of HCl consumed 1 mole of chlorine gas is produced.

total moles of HCl ="\frac{212}{36.5}"=5.8


moles of chlorine gas produced = "\frac{5.8}{4}=1.45"

grams of chlorine produced =1.45"\times"71=102.95


As the percent yield is 91 here so finally amount of chlorine gas obtained ="\frac{91}{100}\times102.95=93.7 g"